OpenStudy (anonymous):

how do i find the value of the infinite series for x in the interval -1

OpenStudy (anonymous):

Can someone help please??

OpenStudy (anonymous):

We know that sum ((x^n),n,0,infinity) = 1/(1-x), so sum(2(-1^n)(x^n),n,0,infinity) = 2 sum((-1^n)(x^n),n,0,infinity) = 2 sum((-x)^n,n,0,infinity) = 2 [1/(1-(-x))] = 2/(1+x)